\begin{thebibliography}{99}
   \bibitem{AB}
 C. Aistleitner and  S. Baker.	
 On the pair correlations of powers of real numbers.
 arXiv:1910.01437 [math.NT]
 	
\bibitem{ALL}
 	 C. Aistleitner, G. Larcher and M. Lewko. Additive energy and the Hausdorff dimension of the exceptional set in metric pair correlation problems. With an appendix by Jean Bourgain. Israel J. Math. 222 (2017), no. 1, 463--485.

\bibitem{BCC}
V. Becher, O. Carton and I. Cunningham.
Low discrepancy sequences failing Poissonian pair correlations. Arch. Math. 113 (2019), no. 2, 169--178.



\bibitem{BCGW} 
 T.F. Bloom, S. Chow, A. Gafni and A. Walker. 
 Additive energy and the metric Poissonian property. Mathematika 64 (2018), no. 3, 679--700.
 	
  \bibitem{CLZ}
   S. Chaubey, M. Lanius and A. Zaharescu. 
   Pair correlation of fractional parts derived from rational valued sequences. J. Number Theory 151 (2015), 147--158.
   
   \bibitem{EMV} 
D. El-Baz, J. Marklof and I. Vinogradov. 
The two-point correlation function of the fractional parts of $\sqrt{n}$ is Poisson.
Proc. Amer. Math. Soc. 143 (7), 2815--2828.

  \bibitem{KN}
   L. Kuipers and H. Niederreiter. Uniform distribution of sequences. Pure and Applied Mathematics. Wiley-Interscience [John Wiley \& Sons], New York-London-Sydney, 1974.
   
\bibitem{LT}
T. Lachmann and N. Technau. On exceptional sets in the metric Poissonian pair correlations problem. Monatsh. Math. 189 (2019), no. 1, 137--156.


\bibitem{LarcherStockinger}
 G. Larcher and W. Stockinger.
 Some negative results related to Poissonian pair correlation problems
 To appear in: Disc. Math., 2019. doi.org/10.1016/j.disc.2019.111656
 
 \bibitem{PS}
 I. Pirsic and W. Stockinger. The Champernowne constant is not Poissonian. Funct. Approx. Comment. Math. 60 (2019), no. 2, 253--262. 
 

\bibitem{RudnickSarnak}
  Z. Rudnick and P. Sarnak. The Pair Correlation Function of 
  Fractional Parts of Polynomials. Comm. Math. Phys. 194, (1998). 61 -- 70 .
  
  \bibitem{RZ pc}
Z. Rudnick and A. Zaharescu. A metric result on the pair correlation of fractional parts of sequences, Acta Arith. 89 (1999), 283--293.

\bibitem{Walker}
 A. Walker. The primes are not metric Poissonian. Mathematika 64 (2018), no. 1, 230--236.

\end{thebibliography}

\documentclass[11pt]{amsart}
\usepackage{verbatim}
\usepackage{amssymb, amsmath}
\usepackage{graphicx}
\usepackage{color}

\begin{document}
\newtheorem{thm}{Theorem}[section]
\newtheorem*{thm*}{Theorem}
\newtheorem{lem}[thm]{Lemma}
\newtheorem{prop}[thm]{Proposition}
\newtheorem{cor}[thm]{Corollary}
\newtheorem*{conj}{Conjecture}
\newtheorem{proj}[thm]{Project}

\theoremstyle{definition}
\newtheorem*{defn}{Definition}
\newtheorem*{remark}{Remark}
\newtheorem{exercise}{Exercise}
\newtheorem*{exercise*}{Exercise}

\numberwithin{equation}{section}

\newcommand{\rad}{\operatorname{rad}}

\def\xo{x_{1}}
\def\yo{x_{2}}
\def\xt{x_{3}}
\def\yt{x_{4}}

\def\no{n_{1}}
\def\nt{n_{2}}

\def\etao{\eta}
\def\etat{2\eta}
\def\vx{\mathbf{x}}

\global\long\def\xo{\mathrm{x_{1}}}%
\global\long\def\xt{\mathrm{x_{3}}}%
\global\long\def\yo{\mathrm{x}_{2}}%
\global\long\def\yt{\mathrm{x_{4}}}%

\global\long\def\xbo{\mathrm{\XBox_{1}}}%
\global\long\def\xbt{\mathrm{\XBox_{3}}}%
\global\long\def\ybo{\mathrm{\XBox}_{2}}%
\global\long\def\ybt{\mathrm{\XBox_{4}}}%

\newcommand{\Z}{{\mathbb Z}} 
\newcommand{\Q}{{\mathbb Q}}
\newcommand{\R}{{\mathbb R}}
\newcommand{\C}{{\mathbb C}}
\newcommand{\N}{{\mathbb N}}
\newcommand{\FF}{{\mathbb F}}
\newcommand{\fq}{\mathbb{F}_q}
\newcommand{\rmk}[1]{\footnote{{\bf Comment:} #1}}

\renewcommand{\mod}{\;\operatorname{mod}}
\newcommand{\ord}{\operatorname{ord}}
\newcommand{\TT}{\mathbb{T}}
\renewcommand{\i}{{\mathrm{i}}}
\renewcommand{\d}{{\mathrm{d}}}
\renewcommand{\^}{\widehat}
\newcommand{\HH}{\mathbb H}
\newcommand{\Vol}{\operatorname{vol}}
\newcommand{\area}{\operatorname{area}}
\newcommand{\tr}{\operatorname{tr}}
\newcommand{\norm}{\mathcal N} 
\newcommand{\intinf}{\int_{-\infty}^\infty}
\newcommand{\ave}[1]{\left\langle#1\right\rangle} 
\newcommand{\Var}{\operatorname{Var}}
\newcommand{\Prob}{\operatorname{Prob}}
\newcommand{\sym}{\operatorname{Sym}}
\newcommand{\disc}{\operatorname{disc}}
\newcommand{\CA}{{\mathcal C}_A}
\newcommand{\cond}{\operatorname{cond}} 
\newcommand{\lcm}{\operatorname{lcm}}
\newcommand{\Kl}{\operatorname{Kl}} 
\newcommand{\leg}[2]{\left( \frac{#1}{#2} \right)}  
\newcommand{\Li}{\operatorname{Li}}

\newcommand{\sumstar}{\sideset \and^{*} \to \sum}

\newcommand{\LL}{\mathcal L} 
\newcommand{\sumf}{\sum^\flat}
\newcommand{\Hgev}{\mathcal H_{2g+2,q}}
\newcommand{\USp}{\operatorname{USp}}
\newcommand{\conv}{*}
\newcommand{\dist} {\operatorname{dist}}
\newcommand{\CF}{c_0} 
\newcommand{\kerp}{\mathcal K}

\newcommand{\Cov}{\operatorname{cov}}
\newcommand{\Sym}{\operatorname{Sym}}

\newcommand{\Ht}{\operatorname{Ht}}

\newcommand{\E}{\operatorname{E}} 
\newcommand{\sign}{\operatorname{sign}} 
\newcommand{\meas}{\operatorname{meas}} 

\newcommand{\divid}{d} 

\newcommand{\GL}{\operatorname{GL}}
\newcommand{\SL}{\operatorname{SL}}
\newcommand{\re}{\operatorname{Re}}
\newcommand{\im}{\operatorname{Im}}
\newcommand{\res}{\operatorname{Res}}
\newcommand{\Str}{\mathcal S^{\rm trunc}} 
 \newcommand{\length}{\operatorname{length}}
 
\newcounter{ToDo}
\newcommand{\todo}[1]{\stepcounter{ToDo}%
	({\color{blue}TODO~\arabic{ToDo}: {#1}})%
}

\title[Pair correlation of lacunary sequences]{The metric theory of the pair correlation function of real-valued lacunary sequences}
\author{Ze\'ev Rudnick and Niclas Technau}
\date{\today}
 \address{School of Mathematical Sciences, Tel Aviv University, Tel Aviv 69978, Israel}
 
\email{rudnick@tauex.tau.ac.il}
\email{niclast@mail.tau.ac.il}
\thanks{This result is part of a project that received funding from the European Research Council (ERC) under the European Union's  Horizon 2020 research and innovation programme  (Grant agreement No.    786758).}
\subjclass[2010]{11J54; 11J71}

\keywords{Pair correlation; Poisson statistics; lacunary sequence}
\begin{abstract}
Let $\{ a(x) \}_{x=1}^{\infty}$ be a positive, real-valued, 
lacunary sequence.
This note shows that the pair correlation function 
of the fractional parts of the dilations $\alpha a(x)$
is Poissonian for Lebesgue almost every $\alpha\in \mathbb{R}$. 
By using harmonic analysis, our result 
--- irrespective of the choice of the real-valued sequence 
$\{ a(x) \}_{x=1}^{\infty}$ ---
can essentially be reduced to showing that
the number of solutions to the Diophantine inequality 
$$
\vert n_1 (a(x_1)-a(y_1))- n_2(a(x_2)-a(y_2)) \vert < 1 
$$
in integer six-tuples $(n_1,n_2,x_1,x_2,y_1,y_2)$ located in the box 
$[-N,N]^6$ with the ``excluded diagonals'', that is 
$$x_1\neq y_1, \quad x_2 \neq y_2, \quad (n_1,n_2)\neq (0,0),$$ 
is at most $N^{4-\delta}$ for some fixed $\delta>0$, 
for all sufficiently large $N$. 

\end{abstract}
\maketitle

\section{Introduction}
A sequence of points $\{\theta_n\}_{n=1}^\infty$  
is uniformly distributed modulo one if given any fixed interval  $I$ in the unit circle $\R/\Z$, 
 the proportion of fractional parts $\theta_n \bmod 1$ which lie in $I$ tends to the length of the interval $I$, that is 
 \begin{equation*}
 \#\{n\leq N: \theta_n\bmod 1 \in I\} \sim  \length(I) \cdot N, \quad N\to \infty .
 \end{equation*}
 
 We study the {\em pair correlation function}   $R_2$, defined for every fixed interval $I\subset \R$ by the property that
\[
\lim_{N\to \infty} \frac 1N \#\{ 1\leq m\neq n\leq N: |\theta_n-\theta_m| \in \frac 1N I \} = \int_I R_2(x)dx
 \]
assuming that the limit exists. For a random sequence of $N$ elements, that is $N$ uniform independent random variables in $[0,1)$ (the Poisson model), the limiting pair correlation function is  $R_2(x)\equiv 1$.

There are very few positive results on the pair correlation function available for specific sequences,  a notable exception being the fractional parts of $\sqrt{n}$ \cite{EMV}; 
a more tractable problem is to randomize (a ``metric'' theory, in the terminology of uniform distribution theory) by looking at 
random multiples\footnote{A different notion of randomizing has recently been investigated in \cite{AB}, which studies the pair correlation of the sequence $\alpha^n \bmod 1$ with $\alpha$ random.}
$\theta_n=\alpha a(n) \bmod 1$, for almost all $\alpha$. 
 There is a well-developed metric theory of the pair correlation function for  {\em integer} valued sequences $\{a(n)\}_{n=1}^\infty$, initiated in  \cite{RudnickSarnak}, where polynomial sequences such as $a(n)=n^d$, $d\geq 2$, are studied,  
 with several developments in the last few years, see e.g. \cite{ALL, BCC, CLZ, LT, LarcherStockinger, PS, RZ pc, Walker}.  
 In this note we study the case of real-valued lacunary sequences:  
Let $a(x)>0$ be a lacunary sequence of positive reals, that is there is some $C>1 $ so that for all integers $x\geq 1$, 
\[
a(x+1)\geq C a(x).
\] 
For instance, we can take $a(x) = e^x$. It is known  that for almost all $\alpha$, the sequence $\alpha a(x)\bmod 1$ is uniformly distributed mod one \cite[Chapter 1, Corollary 4.3]{KN}. 
Here and throughout this note ``almost all'' is meant
with respect to the Lebesgue measure on $\R$.

\begin{thm}\label{thm:lacunary} 
Assume that $\{a(x)\}_{x=1}^\infty$
is a lacunary sequence of positive reals. 
Then the pair correlation function of 
the sequence $\{\alpha a(x)\}_{x=1}^\infty$ 
is Poissonian for almost all $\alpha$. 
\end{thm}

When $a(x)$ takes {\em integer} values,   \cite{RZ pc}  showed that for almost all $\alpha$, the pair correlation function is Poissonian. 
The case of  pair correlation of sequences of {\em rationals} $x_n=a_n/b_n$ with $a_n$ integer-valued and lacunary and $b_n$ integer-valued and (roughly speaking) sufficiently small (e.g.  $a_n/b_n=2014^n/[\log\log n]$) was treated in \cite{CLZ}.   Here we treat any real-valued sequences.

We will reduce the problem to giving a bound for the number of lattice points satisfying a Diophantine inequality: 
Let 
$$M=N^{1+\varepsilon}, \quad K=N^\varepsilon
$$ 
 and assume that there is some $\delta>0$ so that for all $\varepsilon>0$ sufficiently small
\begin{equation}\label{cond expec} \tag{A}
\#\{1\leq n\leq M, 1\leq x\neq y\leq N: n|a(x)-a(y)|<K \} 
\ll N^{2-\delta}.
\end{equation}
Let $\mathcal S(N)$ be the set of integer six-tuples with 
\[
1\leq y_i\neq x_i\leq N, \quad 1\leq |n_i|\leq M, \quad (i=1,2),
\]
satisfying 
\[
| n_1(a(x_1)-a(y_1))- n_2(a(x_2)-a(y_2)) |< K.
\]
Assume that 
\begin{equation}\label{lacunary ineq}\tag{B}
\#\mathcal S(N) \ll  N^{4-\delta}.
\end{equation}

\begin{thm}\label{prop:Reduce pc to ineq}
Let $\{a(x)\}_{x=1}^\infty$ be a sequence of distinct 
positive reals. 
Assume that  \eqref{cond expec} and \eqref{lacunary ineq} hold 
for some $\delta>0$.    
Then the pair correlation function of $\alpha a(x)$ 
is Poissonian 
for Lebesgue almost all $\alpha\in \mathbb{R}$.
\end{thm}

In the case of integer-valued sequences, the almost sure convergence of the pair correlation function to the Poisson limit (metric Poisson pair correlation) follows \cite{RudnickSarnak, RZ pc} 
from  a similar bound for the equation 
$$
n_1 (a(x_1)-a(y_1))- n_2(a(x_2)-a(y_2))=0
$$ 
See  \cite{ALL, BCGW} for  a  streamlined criterion for metric Poisson pair correlation in terms of the {\em additive energy}  of the sequence. 

In \S~\ref{sec:lacunary}, we verify that that  \eqref{cond expec} and \eqref{lacunary ineq} hold for lacunary sequences, hence obtain Theorem~\ref{thm:lacunary}.

 \section{The pair correlation function}\label{sec:pc} 
 To study the pair correlation function, we use a smooth count cf. \cite{RudnickSarnak}: 
For $f\in C_c^\infty(\R)$ or $f$ being an indicator function
of a compact interval, set 
 \[
 F_N(x) = \sum_{j\in \Z} f\big(N(x+j)\big) 
 \]
 which is periodic and localized on scale $1/N$. 
 For a sequence $\{\theta_n\}_{n=1}^\infty\subset \R/\Z$, 
 we define its pair correlation function by
\begin{equation}
R_2(f,N)(\{\theta_n\}_{n=1}^\infty)= \frac{1}{N} 
\sum_{1\leq m\neq n\leq N} 
F_N(\theta_n-\theta_m).
\end{equation}  
In particular, for a fixed sequence $\{x_n\}_{n=1}^\infty$, 
we take $\theta_n=\alpha x_n\bmod 1$,  
and abbreviate the pair correlation function 
$R_2(f,N)(\{\theta_n\}_{n=1}^\infty)$, having fixed $f$,
by $R_2(f,N)(\alpha)=R_2(\alpha)$. 

It suffices to restrict $\alpha$ to lie in a fixed  finite interval  and to consider a smooth average: 
Let   $\rho\in C_c^\infty(\R)$, $\rho\geq 0$, be a smooth, compactly supported, non-negative weight function, normalized to give a probability density:   $ \int_{\R}\rho(\alpha)d\alpha=1$.  We define a smooth average 
 \begin{equation}\label{eq: smooth ave}
  \ave{X} = \int_{\R} X(\alpha)\rho(\alpha) \, \mathrm{d}\alpha . 
  \end{equation}  

 \subsection{The expected value}
\begin{lem}\label{lem:expectation} 
Let $M=N^{1+\varepsilon}$, $K=N^\varepsilon$ and assume
\eqref{cond expec} holds. 
Then the expected value of $R_2(f,N)(\alpha)$ is 
\[
\ave{R_2(f,N)} =\intinf f(x)dx +O(N^{-\delta}) 
\]
\end{lem}
\begin{proof}
Let $f\in C_c^\infty(\R)$. 
By using Poisson summation, we have the expansion  
 \[
 F_N(x) = \sum_{j\in \Z} f\big(N(x+j)\big) = \frac 1N\sum_{n\in \Z} \widehat f\Big(\frac nN\Big) e(nx)
 \]
 with $e(z):=e^{2\pi i z}$, which gives 
\begin{equation}\label{eq: Fourier expansion of R}
 R_2(\alpha) = \frac 1{N^2}\sum_{n\in \Z} 
 \widehat f\Big(\frac nN\Big) S_{n,N}(\alpha) 
\end{equation} 
where 
$$
S_{n,N}(\alpha) = \sum_{1\leq x \neq y \leq N}
e(\alpha n(a(x)-a(y))).
$$
Therefore the expected value is 
\[
\ave{R_2}=\intinf R_2(\alpha)\rho(\alpha)d\alpha = 
\frac 1{N^2}\sum_{n\in \Z} 
 \widehat f\Big(\frac nN\Big) \sum_{1\leq x \neq y \leq N} 
 \^\rho(n(a(x)-a(y))).
\]
The zero mode $n=0$ gives a contribution of 
\[
\frac 1{N^2}\^f(0) N(N-1) = \intinf f(x)dx (1+O(1/N)). 
\]

We split the sum over non-zero modes into two terms: Those with $1\leq |n|\leq M=N^{1+\varepsilon}$, and those with $|n|>M$. 
To treat the contribution of modes with $|n|>M=N^{1+\varepsilon}$, we use $|\^f(x)| \ll x^{-A}$  and $|\^\rho|\ll 1$ to  bound that term by 
\[
 \frac 1{N^2} \sum_{|n|>M} \Big(\frac nN \Big)^{-A} \sum_{1\leq x\neq y\leq N} 1 = \frac {N^A}{M^{A-1}} \ll \frac{1}{N^{1-\varepsilon}}
\]
on choosing $A=2/\varepsilon$. 

To bound the contribution of modes with $1\leq |n|\leq M$, we separate into a contribution of terms with $|n(a(x)-a(y))|<K$ and the rest. 

We use $|\^\rho|, |\^f|\ll 1$ to obtain that the contribution of terms with $|n(a(x)-a(y))|<K$ is 
\[
\ll \frac 1{N^2}\#\{1\leq n<N^{1+\varepsilon}, 1\leq y\neq x\leq N: n(a(x)-a(y))<N^\varepsilon\}.
\]
By \eqref{cond expec}, this is $\ll N^{-\delta}$. 

The contribution of  terms with $|n(a(x)-a(y))|>K$  is bounded using 
$$|\^\rho(n(a(x)-a(y)))\ll |n(a(x)-a(y))|^{-A}\leq K^{-A} =N^{-2}$$
and $|\^f|\ll 1$ by 
\[
\frac 1{N^2}\sum_{\substack{1\leq |n|\leq M\\1\leq x\neq y\leq N\\  |n(a(x)-a(y))|>K}} |\^f(\frac nN)|  
\^\rho( n(a(x)-a(y)) |
\ll \frac 1{N^2}\sum_{\substack{1\leq |n|\leq M\\1\leq x\neq y\leq N}} \frac 1{N^2} \leq \frac{M}{N^2}
\]
which is $\ll N^{-1+\varepsilon}$. 
\end{proof}

\subsection{The variance}
\begin{prop} \label{prop:lacunary variance} 
Assume that $a(x)$ is a sequence of   real numbers such that \eqref{cond expec} and \eqref{lacunary ineq} hold. Then 
\[
\ave{\Big| R_2(f,N)-\intinf f(x)dx \Big|^2}\ll N^{-\delta}.
\]
 \end{prop}
\begin{proof}
By Cauchy-Schwarz, 
\begin{align*}
\ave{\Big| R_2(f,N)-\intinf f(x)dx \Big|^2} & \leq 
 2\ave{\Big| R_2(f,N)-\ave{R_2} \Big|^2}
\\
& +2 \ave{\Big| \ave{R_2}-\intinf f(x)dx \Big|^2}.
\end{align*}
By Lemma~\ref{lem:expectation}, 
$$ \ave{\Big| \ave{R_2}-\intinf f(x)dx \Big|^2}\ll N^{-2\delta}.
$$ 
We now show that 
\begin{equation}\label{bound on var}
 \Var R_2=\ave{\Big| R_2(f,N)-\ave{R_2}\Big|^2} \ll N^{-\delta} 
 \end{equation}
which will  prove Proposition~\ref{prop:lacunary variance}.

 To prove \eqref{bound on var}, it suffices to show by \eqref{lacunary ineq}  that
 \begin{equation} \label{lem: variance bounded in terms of counting}
\Var R_2 \ll_f \frac{\#\mathcal S(N) }{N^4}
\end{equation}

By using the expansion \eqref{eq: Fourier expansion of R}, the variance can be written as
  \begin{equation}\label{eq: variance in terms of Fourier coefficients}
   \Var(R_{2})=\frac{1}{N^{4}} \sum_{(n_1,n_2)  \in \mathbb{Z}^{2}\setminus\{0\}} 
  \widehat{f}\Big(\frac{n_1}{N}\Big) \widehat{f}\Big(\frac{n_2}{N}\Big)
  w(\no,\nt,N).
  \end{equation}
where  for integers $n_1,n_2$, we let
$$w(\no,\nt, N) = 
\sum_{\substack{1 \leq x_1 \neq x_3 \leq N, \\ 
1\leq x_2 \neq x_4 \leq N}}
\widehat{\rho}\Big(\no (a(x_3)-a(x_1))
-\nt(a(x_4)-a(x_2))\Big)$$
and $\rho$  as in \eqref{eq: smooth ave}.

Due to the rapid decay  
of $\widehat{f}$, the contribution from the range in which $|\no|$ or $|\nt|$ exceeds $M=N^{1+ \varepsilon}$ 
is negligible, as we will argue now.
We detail only the case  $\max\{|\no|,|\nt|\} =\no \geq M$,
since the other case can be done similarly. 
We observe the trivial bound 
$\vert w(\no,\nt,N) \vert \ll N^4$. Moreover, 
$$n_1=n^{\varepsilon/2}_{1}n^{1-\varepsilon/2}_{1}
\geq n^{\varepsilon/2}_{1} N^{1+\varepsilon/2-\varepsilon^{2}/2}
$$
which, since $\varepsilon$ is small, yields $n_1 > n^{\varepsilon/2}_{1} N^{1+\varepsilon/3}$.
Hence, the contribution to the right hand side of \eqref{eq: variance in terms of Fourier coefficients} 
arising from the terms with $\max\{|\no|,|\nt|\} =\no \geq M=N^{1+ \varepsilon}$ and $n_2 \neq 0$ is
 \begin{align*}
\ll \frac 1{N^4} \sum_{\substack{\no,\nt
 \in \mathbb{Z}\setminus\{0\}\\ 
 \vert \no \vert >N^{1+ \varepsilon}}} 
 \Big(\frac{\no}{N}\Big)^{-  18/\varepsilon}  
 \sum_{n_2\neq 0} \widehat f\Big(\frac{n_2}{N}\Big)
 N^{4}
 \ll \frac 1{N^4}.
 \end{align*}
Moreover, the terms satisfying $\max\{|\no|,|\nt|\} =\no \geq N^{1+ \varepsilon}$, and $n_{2}=0$ 
are in absolute value 
 $$
 \ll \frac 1{N^4}\sum_{\vert \no \vert \geq N^{1+\varepsilon}} 
 \widehat{f}\Big(\frac{\no}{N}\Big) N^{4}\ll \frac 1{N^4}.
 $$
So, the upshot is that on the right hand side of \eqref{eq: variance in terms of Fourier coefficients} 
the sum over all $(n_1,n_2)$ with $\max(|n_1|,|n_2|)>N^{1+\varepsilon}$
   
contributes $\ll N^{-4}.$
By the rapid decay  
of $\widehat{\rho}$, we can dispose of the regime where 
 $ \vert n_1 (a(x_3)-a(x_1))-n_2(a(x_4)-a(x_2))  \vert \geq N^{\varepsilon}.
 $

By bounding $\hat{\rho}$ trivially, we find that
\[
\Var R_2 \ll \frac{\#\mathcal S(N)}{N^4} + O\Big(\frac 1{N^4}\Big) .
\]
Since $\#\mathcal S(N)\geq N^3$, we obtain \eqref{lem: variance bounded in terms of counting}. 
\end{proof}

    \section{Almost everywhere convergence: Proof of Theorem~\ref{prop:Reduce pc to ineq}}

 We now deduce  almost everywhere convergence   
 from a polynomial variance bound.

 \subsection{Preparations}\label{secion: preparations}
 We need a general property of the pair correlation function. Recall that for any sequence of points 
 $\{\theta_n\}_{n=1}^\infty \subset \R/\Z$, we defined
 \[
 R_2(f,N) = \frac 1N \sum_{1\leq j\neq k\leq N} F_N(\theta_j-\theta_k)
 \]
 with $ F_N(x) = \sum_{j\in \Z} f(N(x-j))$. 
 
 \begin{lem}\label{lem: conv subsequences}
 Suppose there is a strictly increasing sequence 
 $\{ N_m \}_{m=1}^\infty \subseteq \mathbb{Z}_{\geq 1} $, with 
 \[
\lim_{m\to \infty} \frac{N_{m+1}}{N_m} = 1
 \] 
 so that for all $f\in C_c^\infty(\R)$, 
 \begin{equation}\label{convergence on subsequence}
 \lim_{m\to \infty} R_2(f,N_m) = \intinf f(x)dx .
 \end{equation}
 Then we can pass from the sub-sequence to the set of all integers:
\begin{equation}\label{convergence on full sequence}
 \lim_{N\to \infty} R_2(f,N) = \intinf f(x)dx
 \end{equation}
 for all $f\in  C_c^\infty(\R)$.
 \end{lem}
 \begin{proof}
We will first deduce that \eqref{convergence on subsequence} holds for the indicator functions 
 \[
 I_s(x) = \begin{cases} 1,& |x|<s/2,\\ 0,&{\rm otherwise}, \end{cases}
 \]
 by approximating with smooth functions, and show that \eqref{convergence on full sequence} holds for the functions $I_s$, and then deduce by approximating a general even smooth $f\in C_c^\infty(\R)$  by linear combinations of $I_s$ 
 that \eqref{convergence on full sequence} holds for all such $f$.
 Note that for odd smooth $f\in C_c^\infty(\R)$, we have 
 $F_N(-x)=-F_N(x)$ which entails $R_2(f,N)=0$, 
 so the pair correlation function $R_2(f,N)$ 
 converges trivially to the right limit.

 From the definition of $R_2(I_s,N)$ we have a monotonicity property: 
Let $0<\varepsilon<1$.   If $  (1-\varepsilon)N'<N<N'$   and $N''<N<(1+\varepsilon)N''$ then 
 \begin{equation}\label{monotonicity prop}
  (1-\varepsilon) R_2(I_{(1-\varepsilon)s} ,N'') \leq R_2(I_s,N) \leq \frac 1{1-\varepsilon} R_2(\ I_{s/(1-\varepsilon)},N') .
 \end{equation}
    Indeed, using positivity of $ I_s$ (hence of $F_N$)
 \[
 N\cdot R_2(I_s,N) = \sum_{1\leq j\neq k\leq N} F_N(\theta_j-\theta_k) \leq \sum_{1\leq j\neq k\leq N'} F_N(\theta_j-\theta_k)  .
 \]
 Now if   $1>N/N'\geq 1-\varepsilon>0$ then since $I_s$ is even and decreasing on $[0,\infty)$,  we have 
  \[
 I_s(Ny) = I_s\Big(N' y\cdot \frac{N}{N'}\Big) 
 \leq I_s(N'y(1-\varepsilon)) = I_{s/(1-\varepsilon)}(N'y   ).
 \]
So 
 \[ 
 F_N(x) =    \sum_{j\in \Z} I_s(N \cdot(x-j)) 
 \leq \sum_{j\in \Z} I_{s/(1-\varepsilon)} (N'\cdot(x-j)) 
 = \tilde F_{N'}(x) 
\] 
where $\tilde F_{N'}(y) = \sum_{j\in \Z}  I_{s/(1-\varepsilon)}({N'}(y-j))$. Hence 
 \[
 R_2(I_s,N) \leq \frac{N'}{N} R_2(I_{s/(1-\varepsilon)},N') \leq \frac 1{1-\varepsilon}  R_2(I_{s/(1-\varepsilon) },N')
\]
which proves the upper bound in \eqref{monotonicity prop}. 
The lower bound of \eqref{monotonicity prop} follows from switching the roles of $N$ and $N''$ and inserting in the upper bound.

 Next, fix $\varepsilon\in (0,1)$ small, let $N\gg 1$, and   take $m\gg 1$ so that 
 \[
 N_m<N_{m+1}<(1+\varepsilon)N_m
 \]
 and so if $N_m\leq N<N_{m+1}$ then 
 \[
  (1-\varepsilon)N_{m+1} <N < N_{m+1}, \quad N_{m}\leq N<(1+\varepsilon)N_m .
 \]
 Then for all $s>0$, we have
 \[
 (1-\varepsilon) R_2(I_{(1-\varepsilon)s}, N_m) \leq R_2(I_s,N) \leq \frac 1{1-\varepsilon} R_2(I_{s/(1-\varepsilon)},N_{m+1}) .
 \] 
 Taking $m\to \infty$, we find by \eqref{convergence on subsequence}
 \[
 \limsup_{N\to \infty} R_2(I_s,N) \leq \frac 1{1-\varepsilon}  \intinf I_{s/(1-\varepsilon)} dx = \frac{s}{(1-\varepsilon)^2}
  \]
 and 
 \[
 \liminf_{N\to \infty} R_2(I_s,N)  \geq (1-\varepsilon) \intinf I_{ (1-\varepsilon)s} dx  = (1-\varepsilon)^2s .
 \]
 Since $\varepsilon>0$ is arbitrary, we finally obtain 
 \[
 \lim_{N\to \infty} R_2(I_s,N) = s = \intinf I_s(x)dx
 \]
 so that \eqref{convergence on full sequence} holds for all indicator functions $I_s$. 
 Therefore \eqref{convergence on full sequence} holds for all test functions $f\in C_c^\infty(\R)$.
 \end{proof} 
 
 \subsection{Proof of Theorem~\ref{prop:Reduce pc to ineq}}
 
 It suffices to show that for almost every $\alpha$ 
 in a fixed compact interval $I$ we have 
  \begin{equation}\label{full conv}
\lim_{N\to \infty}R_2(f,N)(\alpha)=\intinf f(x)dx
 \end{equation} 
 for all $f\in C_c^\infty(\R)$. 
 
Let $\rho\in C^\infty_c(\R)$ be a non-negative function majorizing the indicator function of the interval $I$: 
$\mathbf 1_{I}\leq \rho$. Then from the variance bound of Proposition~\ref{prop:lacunary variance}, we find that  
 for some $ \delta>0$, for all $f\in C_c^\infty(\R)$, 
 \[
   \int_{I} \Big |R_2(f,N_m)(\alpha)-\intinf f(x)dx \Big|^2 \rho(\alpha)d\alpha \ll_f N^{-\delta} .
 \]
 Hence for the sequence
 \[
 N_m = \lfloor m^{2/\delta} \rfloor
 \]
 we have that for almost all $\alpha \in I$,
 \begin{equation}\label{subsequential conv}
\lim_{m\to \infty} R_2(f,N_m)(\alpha) = \intinf f(x)dx
 \end{equation}
 for all $f$. Indeed, for each fixed $f$  set 
 \[
 X_m (\alpha) = \Big|R_2(f,N_m)(\alpha)-\intinf f(x)dx \Big|^2 .
 \]
Then
 \[
 \int_{I} X_m(\alpha)d\alpha \leq  \int_{-\infty}^{\infty}   X_m (\alpha)  \rho(\alpha)d\alpha  \ll   \frac 1{N_m^{\delta}} \ll \frac 1{m^2} .
 \] 
 Therefore 
  \[
 \int_{I} \Big(  \sum_{m\geq 1} X_m(\alpha) \Big) d\alpha  \leq \sum_{m\geq 1} \int_{-\infty}^\infty X_m(\alpha) d\alpha \ll_f   \sum_{m\geq 1} \frac 1{m^2}<\infty
 \]
 so that $\sum_{m\geq 1} X_m(\alpha)$ converges 
 for almost all $\alpha\in I$. Thus 
 $$\lim_{m\to \infty}X_m(\alpha)=0$$ 
 for almost all $\alpha$, i.e. \eqref{subsequential conv} holds for our specific $f$ for almost all $\alpha\in I$. 
 
 By a diagonalization argument (see \cite{RudnickSarnak})  there is a set of $\alpha$ whose complement has measure zero so that \eqref{subsequential conv} holds for all $f$. Since $N_{m+1}/N_m\to 1$, 
 we can use Lemma~\ref{lem: conv subsequences} to deduce \eqref{full conv} holds, proving Theorem~\ref{prop:Reduce pc to ineq}.

 \section{Lacunary sequences}\label{sec:lacunary}
 
 From now on, we assume that  $\{a(x)\}_{x=1}^\infty$ 
 is a lacunary sequence of (strictly) positive reals, 
 that is there is some $C>1 $ 
so that for all integers $x\geq 1$, 
\[
a(x+1)\geq C a(x)
\] 
for all $x\geq 1$. Consequently, we have for all $x\geq y\geq 1$ that
\[
a(x)\geq C^{x-y} a(y) .
\]
We will show that \eqref{cond expec} and \eqref{lacunary ineq} hold, hence proving that the pair correlation function of 
$\{\alpha a(x) \mod 1\}_{x=1}^\infty$ is Poissonian for almost all $\alpha$, that is Theorem~\ref{thm:lacunary}.

 \subsection{The condition  \eqref{cond expec}}
 \begin{lem}
Assume that the sequence $\{a(x)\}_{x=1}^\infty$ is lacunary. Then \eqref{cond expec} holds, in fact with a bound of 
\[
\#\{1\leq n\leq M, 1\leq x\neq y\leq N: n|a(x)-a(y)|<K \} 
\ll N^{2\varepsilon}.
\]
\end{lem}
\begin{proof}
Since the sequence is lacunary, we have for $y<x$ that 
\[
a(x)-a(y) \geq a(x)\Big(1-\frac 1{C^{x-y}}\Big)\geq C^x\Big(1-\frac 1{C^{x-y}}\Big)
\]
which is $\gg N^\varepsilon$ as soon as  $x\geq \epsilon \log_C N$
where $\log_C N=(\log N)/\log C$ . Hence to satisfy the inequality we must have $n<N^\varepsilon$, and $y<x\ll \log N$, so that we have at most $N^{2\varepsilon}$ solutions. 

\end{proof}

 \subsection{The condition \eqref{lacunary ineq}}
 
\begin{prop}\label{prop:counting}
Assume that  $\{a(x)\}_{x=1}^\infty$ is a lacunary sequence of positive real numbers, and that $N^\gamma\ll M\ll N^\Gamma$ for some $0<\gamma<\Gamma<2$. Then 
\[
\#\mathcal S(N)   \ll MN^2(\log M)^2.
\]
\end{prop}
In view of Theorem~\ref{prop:Reduce pc to ineq}, we deduce Theorem~\ref{thm:lacunary}. 

\begin{proof}
The proof is a  modification of  \cite[Proposition 2]{RZ pc}: We are given the inequality
\begin{equation}\label{lacunary ineq2}
| n_1 (a(x_1)-a(y_1))- n_2(a(x_2)-a(y_2)) | < K.
\end{equation}
We may assume that $n_i>0$, and $1\leq y_i<x_i\leq N$, $i=1,2$, and that $x_1\geq x_2$. In particular, we may then assume that $x_1\geq 4\log_C M \gg \log N$, 
because the number of such tuples with $x_1\ll \log N$ is at most 
$O(M^2(\log N)^4)$, which  is admissible 
(that is, $o(MN^2(\log N)^2)$) if $M=O(N^\Gamma)$ for $\Gamma<2$.

We fix $n_1,x_1,y_1$, and first show that (recall $x_1\geq x_2$) 
\begin{equation}\label{diff between s_i}
x_1 - x_2 \leq 2\log_C M .
\end{equation}
Indeed,  we have a lower bound 
\[
  n_1(a(x_1)-a(y_1)) \geq 1\cdot (a(x_1)-a(x_1-1)) 
  \geq \Big(1-\frac 1C \Big) a(x_1)
\]
(since $y\leq x_1-1$), and an upper bound 
\[
n_2(a(x_2)-a(y_2))   \leq M a(x_2) =a(x_1)M \frac{a(x_2)}{a(x_2+(x_1-x_2))} \leq \frac{M}{C^{x_1-x_2}}a(x_1) 
\]
since $a(x+h)\geq C^h a(x)$ for $h\geq 1$. Hence
\[
n_1(a(x_1)-a(y_1))
- n_2(a(x_2)-a(y_2)) 
\geq \Big(1-\frac 1C\Big) a(x_1) - \frac{M}{C^{x_1-x_2}}a(x_1) 
\]
Assuming that
$ x_1-x_2>2\log_C M $
gives in particular 
\[
1-\frac 1C -\frac{M}{C^{x_1-x_2}} >1-\frac 1C-\frac 1M 
> \frac 12 \Big(1-\frac 1C\Big)>0
\]
for sufficiently large $N$.
The condition \eqref{lacunary ineq2}   now forces
\[
\frac 12(1-\frac 1C)<  (1-\frac 1C)   - \frac{M}{C^{x_1-x_2}}  \leq \frac K{a(x_1)} \ll \frac K{C^{x_1}}
\]
which forces $x_1 \ll \log_C K\leq \varepsilon \log M$, which we assumed was not the case. Thus we may assume that $x_1-x_2>2\log_C M$, which forces 
$x_2\geq 2\log_CM$ 
since $x_1>4\log_C M$. 

Now fix $x_2$ as well; 
then $n_2$ will be determined by $y_2$, because 
\[
n_2 = \frac{n_1 (a(x_1)-a(y_1))}{a(x_2)-a(y_2)} + O\Big(\frac K{ a(x_2)-a(y_2)}\Big)
\]
and since $a(y)$ is lacunary, $K/(a(x_2)-a(y_2)) =o(1)$ 
if $ y_2 \geq \log_C N$, because 
\[
a(x_2)-a(y_2)  \geq a(x_2)-a(x_2-1)\geq a(x_2)(1-\frac 1C) \gg C^{x_2}\geq M^2
\]
since $x_2>2\log_CM$. 

So we will be done if we show that there is at most one choice of $y_2$ such that  $x_2-y_2>2\log_C M$. Indeed, if there are two pairs $(y_2,n_2)$ and  $(y_2',n_2')$ 
for which \eqref{lacunary ineq2} holds (recall all other variables are now fixed), 
with $x_2-y_2>2\log_C M$, $x_2-y_2'>2\log_C M$, then since
\[
a(y_2) \leq \frac{a(x_2)}{C^{x_2-y_2}}\leq \frac{a(x_2)}{M^2}
\]
we find that   \eqref{lacunary ineq2} implies 
\[
\begin{split}
n_1(a(x_1)-a(y_1))  &= n_2(a(x_2)-a(y_2)) +O(K) 
\\
&= n_2 a(x_2)\Big( 1+ \frac{a(y_2)}{a(x_2)} 
+O\Big(\frac K{n_2a(x_2)}\Big)\Big) 
\\
&= n_2 a(x_2)\Big( 1+ O\Big(\frac{K}{M^2}\Big)\Big)
\end{split}
\]
since $n_2a(x_2)\geq a(x_2)\geq C^{x_2} \gg M^2$ if $x_2\geq 2\log_C M$, and $a(x_2)/a(y_2)\geq C^{x_2-y_2}\gg M^2$.  
If $n_2',y_2'$ is another such pair then we also find
\[
n_1(a(x_1)-a(y_1)) = n'_2 a(x_2)
\Big( 1+ O \Big(\frac{K}{M^2} \Big)\Big)
\]
so that 
\[
n_2 a(x_2)\Big( 1+ O\Big( \frac{K}{M^2} \Big)\Big)
= n'_2 a(x_2)\Big( 1+ O\Big( \frac{K}{M^2}\Big)\Big)
\]
which gives
\[
n_2'=n_2 \Big( 1+ O\Big(\frac{K}{M^2}\Big)\Big) 
= n_2 + O\Big( \frac{K}{M}\Big)
\]
since $n_2\leq M$. Thus for $M\gg N^\gamma$, 
while $K\ll N^\varepsilon = o(M)$, we obtain $n_2'=n_2$.
\end{proof}

\end{document}